# Testing Non-Standard Neutrinos in Purely Leptonic Lepton Decays

### v1

@@ Abstract
We propose a novel method to probe sterile neutrinos using polarization observables in the purely leptonic decays $\ell^{\prime-}\to\ell^{-}\bar{\nu}_{\ell}\nu_{\ell^{\prime}}$. By analyzing angular distributions and asymmetries derived from polarized decay rates, we identify distinctive signatures of sterile neutrino mixing. In particular, we demonstrate that sterile neutrinos can induce singularities in certain asymmetry parameters as functions of the invariant mass squared of the neutrino pair. These singularities occur for sterile neutrino masses $m_{4\nu}$ satisfying $m_{4\nu}^{2}<m_{\ell^{\prime}}^{2}/2$, providing a clear target for experimental investigation. Our results motivate the incorporation of polarized beam sources at future colliders to enhance sensitivity to sterile neutrinos and other new physics.
@@
The observation of neutrino oscillations Davis:1968cp ; Super-Kamiokande:1998kpq ; SNO:2002tuh ; KamLAND:2002uet ; K2K:2002icj , which demonstrate that neutrinos have mass and undergo lepton flavor mixing, stands as one of the most significant discoveries of physics beyond the original Standard Model (SM). Decades of neutrino oscillation experiments have established a consistent framework of neutrino masses and mixing that successfully explains observations from solar, atmospheric, reactor, and accelerator experiments. The neutrino oscillation anomalies observed in these measurements futher suggest the possible existence of additional neutrino states beyond the three well-known neutrino mass eigenstates. Such new states would need to be predominantly electroweak singlets, making them ”sterile” neutrinos Beuthe:2001rc ; Dasgupta:2021ies that would mix with the active neutrinos while interacting only very weakly through SM forces. In the presence of sterile neutrinos, the mixing between flavor eigenstates $\nu_{\ell}$ ($\ell=e,\,\mu,\,\tau$) and the mass eigenstates $\nu_{i}$ is described by a generalized lepton mixing matrix $U_{\ell i}$:
@@
Current studies about sterile neutrinos primarily target neutrion oscillation, neutrinoless double beta decay, branching ratios of leptonic decays. In this work, we investigate, as a complementary signature, angular distribution of polarization observables in purely leptonic decays of leptons, induced by the mixing of a massive sterile neutrino with the active neutrino sector. We analyze the decays $\ell^{\prime-}\to\ell^{-}\bar{\nu}_{\ell}\nu_{\ell^{\prime}}$, where $(\ell^{\prime}=\tau\,,\ell=e,\mu)$ and $(\ell=\mu\,,\ell=e)$, and compute the corrections to SM predictions arising from sterile neutrinos, with a particular focus on their impact on angular distribution of polarization observables. The sterile neutrino mass is considered across over a range from $\sim\!\mathrm{MeV}$ to $\mathrm{GeV}$, with the assumption of $n_{s}=1$ sterile state, denoted $\nu_{4}$; generalization to $n_{s}\geq 2$ is straightforward. A heavy $\nu_{4}$ in this mass range is unstable and therefore avoids the stringent cosmological constraint on the sum of stable neutrino masses, $\sum_{i}m_{\nu_{i}}\lesssim 0.12~\mathrm{eV}$Planck:2018vyg . We restrict our attention to treat $\nu_{4}$ as a Dirac fermion, which automatically satisfy current constraints from neutrinoless double beta decays CUORE:2017tlq ; GERDA:2018pmc .
@@ II Formalism
$\Gamma$ $=$ $\frac{(2\pi)^{4}}{2m_{\ell^{\prime}}}\int\frac{d^{3}p_{\ell}d^{3}p_{i}d^{3}p_{j}}{16\pi^{3}E_{\ell}E_{i}E_{j}}|{\cal M}(\ell^{\prime-}\to\ell^{-}\bar{\nu}_{i}\nu_{j})|^{2}\delta^{4}(p_{\ell}+p_{i}+p_{j}-p_{\ell^{\prime}})=\frac{1}{2^{9}\pi^{5}m_{\ell^{\prime}}}{\cal I}\,.$ (2)
@@
${\cal M_{\lambda\lambda^{\prime}}}(\ell^{\prime-}\to\ell^{-}\bar{\nu}_{i}\nu_{j})$ $=$ $\frac{G_{F}}{\sqrt{2}}U_{\ell j}U^{*}_{\ell^{\prime}i}\bar{u}_{\ell}(\lambda)\gamma^{\mu}(1-\gamma^{5})u_{\ell^{\prime}}(\lambda^{\prime})\bar{u_{i}}\gamma_{\mu}(1-\gamma^{5})\nu_{j}\,.$ (7)
The lepton mixing matrix $U_{\ell j}$ is a $3\times 3$ unitary matrix within the SM, while it is a $4\times 4$ unitary matrix taking into account the stelie neutrino. Note that neutrinos are assumed to be of the Dirac type. Then the squared magnitude is given by
$|{\cal M}(\ell^{\prime-}\to\ell^{-}\bar{\nu}_{i}\nu_{j})|^{2}$ $=$ $\sum_{i,j}\frac{G_{F}^{2}}{2}|U_{\ell j}|^{2}|U^{*}_{\ell^{\prime}i}|^{2}T(\lambda,\lambda^{\prime},m_{i},m_{j})\,.$ (8)
@@
$L^{\mu\nu}$ $=$ $\sum_{\lambda}\bar{u_{i}}\gamma^{\mu}(1-\gamma^{5})\nu_{j}\bar{\nu_{j}}\gamma^{\nu}(1-\gamma^{5})u_{i}$ (10)
@@
$\ell^{\prime\mu}$ $=$ $(m_{\ell^{\prime}},0,0,0)\,$,
$\ell^{\mu}$ $=$ $(p_{\ell}^{0},|\vec{Y}|\sin\theta_{\ell}\cos\phi_{\ell},|\vec{Y}|\sin\theta_{\ell}\sin\phi_{\ell},|\vec{Y}|\cos\theta_{\ell})\,$,
@@
$\zeta_{\ell}^{\dagger}(+)$ $=$ $\left(\begin{array}[]{cc}\cos(\frac{\phi_{\ell}}{2}),&\sin(\frac{\phi_{\ell}}{2})e^{-i\phi_{\ell}}\\ \end{array}\right)\,,\zeta_{\ell}^{\dagger}(-)=\left(\begin{array}[]{cc}-\sin(\frac{\phi_{\ell}}{2})e^{i\phi_{\ell}},&\cos(\frac{\phi_{\ell}}{2})\\ \end{array}\right)\,$, (17)
@@
$p_{i}\cdot\epsilon^{\dagger}(t)$ $=$ $p_{i}^{0},\,\,p_{i}\cdot\epsilon^{\dagger}(0)=-\vec{p_{i}}\cos\theta_{i},\,\,p_{i}\cdot\epsilon^{\dagger}(\pm)=\pm\frac{1}{\sqrt{2}}\vec{p_{i}}\sin\theta_{i}e^{\mp i\phi_{i}}$
$p_{j}\cdot\epsilon^{\dagger}(t)$ $=$ $p_{j}^{0},\,\,p_{j}\cdot\epsilon^{\dagger}(0)=\vec{p_{i}}\cos\theta_{i},\,\,p_{j}\cdot\epsilon^{\dagger}(\pm)=\mp\frac{1}{\sqrt{2}}\vec{p_{i}}\sin\theta_{i}e^{\mp i\phi_{i}}\,.$ (23)
@@
$|F(+,+,t)|^{2}$ $=$ $|F(+,+,0)|^{2}=\frac{m_{\ell^{\prime}}}{m_{\ell}+p^{0}_{\ell}}\frac{(Y^{0}-\vec{p}_{\ell})^{2}}{Y^{2}}(m_{\ell}-\vec{p}_{\ell}+p^{0}_{\ell})^{2}(1+\cos\theta_{\ell})$
$|F(+,+,+)|^{2}$ $=$ $\frac{2m_{\ell^{\prime}}}{m_{\ell}+p^{0}_{\ell}}(m_{\ell}-\vec{p}_{\ell}+p^{0}_{\ell})^{2}(1-\cos\theta_{\ell}),\,\,\,|F(+,+,-)|^{2}=0$
$|F(-,+,t)|^{2}$ $=$ $|F(-,+,0)|^{2}=\frac{m_{\ell^{\prime}}}{m_{\ell}+p^{0}_{\ell}}\frac{(Y^{0}+\vec{p}_{\ell})^{2}}{Y^{2}}(m_{\ell}+\vec{p}_{\ell}+p^{0}_{\ell})^{2}(1-\cos\theta_{\ell})$
$|F(-,+,+)|^{2}$ $=$ $0,\,\,|F(-,+,-)|^{2}=\frac{2m_{\ell^{\prime}}}{m_{\ell}+p^{0}_{\ell}}(m_{\ell}+\vec{p}_{\ell}+p^{0}_{\ell})^{2}(1+\cos\theta_{\ell})\,$, (26)
@@
$w_{++}$ $=$ $64\pi\frac{m_{\ell^{\prime}}}{Y^{2}}[(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}-(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})\vec{Y}](2p_{i}^{0}p_{j}^{0}-\frac{2}{3}|\vec{p}_{i}|^{2})(1+\cos\theta_{\ell})$
$+128\pi m_{\ell^{\prime}}(p^{0}_{\ell}-\vec{Y})(p_{i}\cdot p_{j}-\frac{2}{3}|\vec{p}_{i}|^{2})(1-\cos\theta_{\ell})\,$,
$w_{-+}$ $=$ $64\pi\frac{m_{\ell^{\prime}}}{Y^{2}}[(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})\vec{Y}](2p_{i}^{0}p_{j}^{0}-\frac{2}{3}|\vec{p}_{i}|^{2})(1-\cos\theta_{\ell})\,$, (27)
$+128\pi m_{\ell^{\prime}}(p^{0}_{\ell}+\vec{Y})(p_{i}\cdot p_{j}-\frac{2}{3}|\vec{p}_{i}|^{2})(1+\cos\theta_{\ell})\,$,
@@
$\int_{D_{1}}d\cos\theta_{\ell}\frac{d\Gamma_{+,+}}{dY^{2}d\cos\theta_{\ell}}$ $\equiv$ $A_{1},\,\,\int_{D_{2}}d\cos\theta_{\ell}\frac{d\Gamma_{+,+}}{dY^{2}d\cos\theta_{\ell}}\equiv B_{1}\,$, (29)
@@
In the SM, where neutrino masses can be neglected, the mixing matrix in Eq. (24) is a $3\times 3$ matrix and the polarized differential decay rates are given by
$\frac{d\Gamma^{S}_{+,+}}{dY^{2}d\cos\theta_{\ell}}$ $=$ $\frac{G^{2}_{F}}{2^{5}\pi^{4}m_{\ell^{\prime}}}|\vec{Y}|\left\{[(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}-(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})\vec{Y}]\frac{1+\cos\theta_{\ell}}{3}\right$.
$+\left.\frac{2}{3}Y^{2}(p^{0}_{\ell}-\vec{Y})(1-\cos\theta_{\ell})\right\}\,$,
$\frac{d\Gamma^{S}_{-,+}}{dY^{2}d\cos\theta_{\ell}}$ $=$ $\frac{G^{2}_{F}}{2^{5}\pi^{4}m_{\ell^{\prime}}}|\vec{Y}|\left\{[(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})\vec{Y}]\frac{1-\cos\theta_{\ell}}{3}\right$. (32)
$+\left.\frac{2}{3}Y^{2}(p^{0}_{\ell}+\vec{Y})(1+\cos\theta_{\ell})\right\}\,.$
@@
$\Upsilon^{S}_{1}$ $=$ $\frac{(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})\vec{Y}+2Y^{2}\vec{Y}}{(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+2Y^{2}p^{0}_{\ell}},\,\,$
$\Upsilon^{S}_{2}$ $=$ $\frac{1}{2}\times\frac{2Y^{2}\vec{Y}-(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})\vec{Y}}{(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+2Y^{2}p^{0}_{\ell}},\,\,$
@@
In the presence of a sterile neutrino, additional mixing marix elements with $i,j=4$ are involved. We derive the corrections of asymmetry parameters relative to the SM predictions from the general decay rate expression in Eq. (24). Note that when both neutrinos are sterile mass eigenstates, i.e. $i=j=4$, the corresponding amplitude involves the product $U_{\ell 4}U^{*}_{\ell^{\prime}4}$, which is doubly suppressed by the small active-sterile mixing and can therefore be safely neglected. Consequently, the net effect by sterile neutrinos reduces to the following expression:
$\Delta\Upsilon_{i}$ $\equiv$ $\Upsilon^{B}_{i}-\Upsilon^{S}_{i}\,$, (34)
@@
$\times\left.[\frac{(1-\frac{m^{4}_{4\nu}}{Y^{4}})(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})\vec{Y}+2(Y^{2}-m^{2}_{4\nu}\frac{2Y^{2}-m^{2}_{4\nu}}{Y^{2}})\vec{Y}}{(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+2Y^{2}p^{0}_{\ell}}]\frac{2|\vec{p}_{i}|}{\sqrt{Y^{2}}}-\Upsilon^{S}_{1}\}\right$.\,,
@@
$\times\left.[\frac{2(Y^{2}-m^{2}_{4\nu}\frac{2Y^{2}-m^{2}_{4\nu}}{Y^{2}})\vec{Y}-(1-\frac{m^{4}_{4\nu}}{Y^{4}})(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})\vec{Y}}{(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+2Y^{2}p^{0}_{\ell}}]\frac{|\vec{p}_{i}|}{\sqrt{Y^{2}}}-\Upsilon^{S}_{2}\}\right$.\,,
@@
The SM provides definite predictions for the asymmetry parameters (Eq. II), which are derived from polarized decay rates. Within the SM framework, the difference between these predictions and experimental measurements is expected to be consistent with zero. Any statistically significant deviation from zero at any values of $Y^{2}$ would therefore suggest new physics beyond the SM, such as the existence of sterile neutrinos.

### v2

@@ Abstract
We propose a method to probe sterile neutrinos using polarization observables in the purely leptonic decays $\ell^{\prime-}\to\ell^{-}\bar{\nu}_{\ell}\nu_{\ell^{\prime}}$. By analyzing angular distributions and asymmetries derived from polarized decay rates, we identify distinctive signatures of sterile neutrino mixing. In particular, we demonstrate that sterile neutrinos can induce singularities in certain asymmetry parameters as functions of the invariant mass squared of the neutrino pair. These singularities occur for sterile neutrino masses $m_{4\nu}$ satisfying $m_{4\nu}^{2}<m_{\ell^{\prime}}^{2}/2$, providing a clear target for experimental investigation. Our results motivate the incorporation of polarized beam sources at future colliders to enhance sensitivity to sterile neutrinos and other new physics.
@@
The observation of neutrino oscillations Davis:1968cp ; Super-Kamiokande:1998kpq ; SNO:2002tuh ; KamLAND:2002uet ; K2K:2002icj , which demonstrate that neutrinos have mass and undergo lepton flavor mixing, stands as one of the most significant discoveries of physics beyond the original Standard Model (SM). Decades of neutrino oscillation experiments have established a consistent framework of neutrino masses and mixing that successfully explains observations from solar, atmospheric, reactor, and accelerator experiments. The neutrino oscillation anomalies observed in these measurements further suggest the possible existence of additional neutrino states beyond the three well-known neutrino mass eigenstates. Such new states would need to be predominantly electroweak singlets, making them ”sterile” neutrinos Marquez:2022bpg ; Beuthe:2001rc ; Dasgupta:2021ies that would mix with the active neutrinos while interacting only very weakly through SM forces. In the presence of sterile neutrinos, the mixing between flavor eigenstates $\nu_{\ell}$ ($\ell=e,\,\mu,\,\tau$) and the mass eigenstates $\nu_{i}$ is described by a generalized lepton mixing matrix $U_{\ell i}$:
@@
Current studies about sterile neutrinos primarily target neutrino oscillation, neutrinoless double beta decay, branching ratios of leptonic decays. In this work, we investigate, as a complementary signature, angular distribution of polarization observables in purely leptonic decays of leptons, induced by the mixing of a massive sterile neutrino with the active neutrino sector. We analyze the decays $\ell^{\prime-}\to\ell^{-}\bar{\nu}_{\ell}\nu_{\ell^{\prime}}$, where $(\ell^{\prime}=\tau\,,\ell=e,\mu)$ and $(\ell=\mu\,,\ell=e)$, and compute the corrections to SM predictions arising from sterile neutrinos, with a particular focus on their impact on angular distribution of polarization observables. The sterile neutrino mass is considered across over a range from $\sim\!\mathrm{MeV}$ to $\mathrm{GeV}$, with the assumption of $n_{s}=1$ sterile state, denoted $\nu_{4}$; generalization to $n_{s}\geq 2$ is straightforward. A heavy $\nu_{4}$ in this mass range is unstable and therefore avoids the stringent cosmological constraint on the sum of stable neutrino masses, $\sum_{i}m_{\nu_{i}}\lesssim 0.12~\mathrm{eV}$Planck:2018vyg . We restrict our attention to treat $\nu_{4}$ as a Dirac fermion, which automatically satisfy current constraints from neutrinoless double beta decays CUORE:2017tlq ; GERDA:2018pmc .
@@ II Formalism
$\Gamma$ $=$ $\frac{(2\pi)^{4}}{2m_{\ell^{\prime}}}\int\frac{d^{3}p_{\ell}d^{3}p_{i}d^{3}p_{j}}{(16\pi^{3})^{3}E_{\ell}E_{i}E_{j}}|{\cal M}(\ell^{\prime-}\to\ell^{-}\bar{\nu}_{i}\nu_{j})|^{2}\delta^{4}(p_{\ell}+p_{i}+p_{j}-p_{\ell^{\prime}})\equiv\frac{1}{2^{9}\pi^{5}m_{\ell^{\prime}}}{\cal I}\,.$ (2)
@@
${\cal M_{\lambda\lambda^{\prime}}}(\ell^{\prime-}\to\ell^{-}\bar{\nu}_{j}\nu_{i})$ $=$ $\frac{G_{F}}{\sqrt{2}}U_{\ell j}U^{*}_{\ell^{\prime}i}\bar{u}_{\ell}(\lambda)\gamma^{\mu}(1-\gamma^{5})u_{\ell^{\prime}}(\lambda^{\prime})\bar{u_{i}}\gamma_{\mu}(1-\gamma^{5})\nu_{j}\,.$ (7)
The lepton mixing matrix $U_{\ell j}$ is a $3\times 3$ unitary matrix within the SM, while it is a $4\times 4$ unitary matrix taking into account the sterile neutrino. Note that neutrinos are assumed to be of the Dirac type. Then the squared magnitude is given by
$|{\cal M}(\ell^{\prime-}\to\ell^{-}\bar{\nu}_{j}\nu_{i})|^{2}$ $=$ $\sum_{i,j}\frac{G_{F}^{2}}{2}|U_{\ell j}|^{2}|U^{*}_{\ell^{\prime}i}|^{2}T(\lambda,\lambda^{\prime},m_{i},m_{j})\,.$ (8)
@@
$L^{\mu\nu}$ $=$ $\sum_{\lambda_{\nu}}\bar{u_{i}}\gamma^{\mu}(1-\gamma^{5})\nu_{j}\bar{\nu_{j}}\gamma^{\nu}(1-\gamma^{5})u_{i}$ (10)
@@
$p^{\mu}_{\ell^{\prime}}$ $=$ $(m_{\ell^{\prime}},0,0,0)\,$,
$p^{\mu}_{\ell}$ $=$ $(p_{\ell}^{0},|\vec{Y}|\sin\theta_{\ell}\cos\phi_{\ell},|\vec{Y}|\sin\theta_{\ell}\sin\phi_{\ell},|\vec{Y}|\cos\theta_{\ell})\,$,
@@
$\zeta_{\ell}^{\dagger}(+)$ $=$ $\left(\begin{array}[]{cc}\cos(\frac{\theta_{\ell}}{2}),&\sin(\frac{\theta_{\ell}}{2})e^{-i\phi_{\ell}}\\ \end{array}\right)\,,\zeta_{\ell}^{\dagger}(-)=\left(\begin{array}[]{cc}-\sin(\frac{\theta_{\ell}}{2})e^{i\phi_{\ell}},&\cos(\frac{\theta_{\ell}}{2})\\ \end{array}\right)\,$, (17)
@@
$p_{i}\cdot\epsilon^{\dagger}(t)$ $=$ $p_{i}^{0},\,\,p_{i}\cdot\epsilon^{\dagger}(0)=-|\vec{p_{i}}|\cos\theta_{i},\,\,p_{i}\cdot\epsilon^{\dagger}(\pm)=\pm\frac{1}{\sqrt{2}}|\vec{p_{i}}|\sin\theta_{i}e^{\mp i\phi_{i}}\,$,
$p_{j}\cdot\epsilon^{\dagger}(t)$ $=$ $p_{j}^{0},\,\,p_{j}\cdot\epsilon^{\dagger}(0)=|\vec{p_{i}}|\cos\theta_{i},\,\,p_{j}\cdot\epsilon^{\dagger}(\pm)=\mp\frac{1}{\sqrt{2}}|\vec{p_{i}}|\sin\theta_{i}e^{\mp i\phi_{i}}\,.$ (23)
@@
$|F(+,+,t)|^{2}$ $=$ $|F(+,+,0)|^{2}=\frac{m_{\ell^{\prime}}}{m_{\ell}+p^{0}_{\ell}}\frac{(Y^{0}-|\vec{p}_{\ell}|)^{2}}{Y^{2}}(m_{\ell}-|\vec{p}_{\ell}|+p^{0}_{\ell})^{2}(1+\cos\theta_{\ell})$
$|F(+,+,+)|^{2}$ $=$ $\frac{2m_{\ell^{\prime}}}{m_{\ell}+p^{0}_{\ell}}(m_{\ell}-|\vec{p}_{\ell}|+p^{0}_{\ell})^{2}(1-\cos\theta_{\ell}),\,\,\,|F(+,+,-)|^{2}=0$
$|F(-,+,t)|^{2}$ $=$ $|F(-,+,0)|^{2}=\frac{m_{\ell^{\prime}}}{m_{\ell}+p^{0}_{\ell}}\frac{(Y^{0}+|\vec{p}_{\ell}|)^{2}}{Y^{2}}(m_{\ell}+|\vec{p}_{\ell}|+p^{0}_{\ell})^{2}(1-\cos\theta_{\ell})$
$|F(-,+,+)|^{2}$ $=$ $0,\,\,|F(-,+,-)|^{2}=\frac{2m_{\ell^{\prime}}}{m_{\ell}+p^{0}_{\ell}}(m_{\ell}+|\vec{p}_{\ell}|+p^{0}_{\ell})^{2}(1+\cos\theta_{\ell})\,$, (26)
@@
$w_{++}$ $=$ $64\pi\frac{m_{\ell^{\prime}}}{Y^{2}}[(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}-(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})|\vec{Y}|](2p_{i}^{0}p_{j}^{0}-\frac{2}{3}|\vec{p}_{i}|^{2})(1+\cos\theta_{\ell})$
$+128\pi m_{\ell^{\prime}}(p^{0}_{\ell}-|\vec{Y}|)(p_{i}\cdot p_{j}-\frac{2}{3}|\vec{p}_{i}|^{2})(1-\cos\theta_{\ell})\,$,
$w_{-+}$ $=$ $64\pi\frac{m_{\ell^{\prime}}}{Y^{2}}[(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})|\vec{Y}|](2p_{i}^{0}p_{j}^{0}-\frac{2}{3}|\vec{p}_{i}|^{2})(1-\cos\theta_{\ell})$ (27)
$+128\pi m_{\ell^{\prime}}(p^{0}_{\ell}+|\vec{Y}|)(p_{i}\cdot p_{j}-\frac{2}{3}|\vec{p}_{i}|^{2})(1+\cos\theta_{\ell})\,$,
@@
$\int_{D_{1}}d\cos\theta_{\ell}\frac{d\Gamma_{+,+}}{dY^{2}d\cos\theta_{\ell}}$ $\equiv$ $A_{1},\,\,\int_{D_{2}}d\cos\theta_{\ell}\frac{d\Gamma_{+,+}}{dY^{2}d\cos\theta_{\ell}}\equiv B_{1}\,$,
@@
In the SM ($i,j=1,2,3$), where neutrino masses can be neglected, it follows from Eq. (II) that $|p_{i}|=p_{i}^{0}=|p_{j}|=p_{j}^{0}=\sqrt{Y^{2}}/2$ and $p_{i}\cdot p_{j}=2|p_{i}||p_{j}|=Y^{2}$. Consequently, $w_{\lambda\lambda^{\prime}}$ in Eq. (24) becomes independent of $i$ and $j$. Using the unitarity of the mixing matrix, $\sum_{j}|U_{lj}|^{2}=\sum_{i}|U^{*}_{l^{\prime}i}|^{2}=1$, the polarized differential decay rates are given by
$\frac{d\Gamma^{S}_{+,+}}{dY^{2}d\cos\theta_{\ell}}$ $=$ $\frac{G^{2}_{F}}{2^{5}\pi^{4}m_{\ell^{\prime}}}|\vec{Y}|\left\{[(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}-(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})|\vec{Y}|]\frac{1+\cos\theta_{\ell}}{3}\right$.
$+\left.\frac{2}{3}Y^{2}(p^{0}_{\ell}-|\vec{Y}|)(1-\cos\theta_{\ell})\right\}\,$,
$\frac{d\Gamma^{S}_{-,+}}{dY^{2}d\cos\theta_{\ell}}$ $=$ $\frac{G^{2}_{F}}{2^{5}\pi^{4}m_{\ell^{\prime}}}|\vec{Y}|\left\{[(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})|\vec{Y}|]\frac{1-\cos\theta_{\ell}}{3}\right$. (31)
$+\left.\frac{2}{3}Y^{2}(p^{0}_{\ell}+|\vec{Y}|)(1+\cos\theta_{\ell})\right\}\,.$
@@
$\Upsilon^{S}_{1}$ $=$ $\frac{(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})|\vec{Y}|+2Y^{2}|\vec{Y}|}{(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+2Y^{2}p^{0}_{\ell}},\,\,$
$\Upsilon^{S}_{2}$ $=$ $\frac{1}{2}\times\frac{2Y^{2}|\vec{Y}|-(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})|\vec{Y}|}{(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+2Y^{2}p^{0}_{\ell}},\,\,$
@@
In the presence of a sterile neutrino, one can derive asymmetry parameters $\Upsilon^{B}_{1,2,3}$ having both the SM and new physics contributions from the general decay rate expression in Eq. (24) with additional mixing matrix elements $i,j=4$ involved. The amplitude corresponding to the case where both neutrinos are sterile mass eigenstates ($i=j=4$) involves the product $U_{\ell 4}U^{*}_{\ell^{\prime}4}$, which is doubly suppressed by the small active-sterile mixing and can therefore be safely neglected. In the following discussion, we focus on the new physics contribution in which one of the neutrinos is sterile while the other is a SM neutrino. As a result, the net effect of sterile neutrinos reduces to the expression below:
$\Delta\Upsilon_{1,2,3}$ $\equiv$ $\Upsilon^{B}_{1,2,3}-\Upsilon^{S}_{1,2,3}\,$, (33)
@@
$\times\left.[\frac{(1-\frac{m^{4}_{4\nu}}{Y^{4}})(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})|\vec{Y}|+2(Y^{2}-m^{2}_{4\nu}\frac{2Y^{2}-m^{2}_{4\nu}}{Y^{2}})|\vec{Y}|}{(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+2Y^{2}p^{0}_{\ell}}]\frac{2|\vec{p}_{i}|}{\sqrt{Y^{2}}}-\Upsilon^{S}_{1}\}\right$.\,,
@@
$\times\left.[\frac{2(Y^{2}-m^{2}_{4\nu}\frac{2Y^{2}-m^{2}_{4\nu}}{Y^{2}})|\vec{Y}|-(1-\frac{m^{4}_{4\nu}}{Y^{4}})(m^{2}_{\ell^{\prime}}-m^{2}_{\ell})|\vec{Y}|}{(m^{2}_{\ell^{\prime}}+m^{2}_{\ell})p^{0}_{\ell}-2m_{\ell^{\prime}}m^{2}_{\ell}+2Y^{2}p^{0}_{\ell}}]\frac{|\vec{p}_{i}|}{\sqrt{Y^{2}}}-\Upsilon^{S}_{2}\}\right$.\,,
@@
The SM provides definite predictions for the asymmetry parameters (Eq. II), which are derived from polarized decay rates. Within the SM framework, the difference between these predictions and experimental measurements is expected to be consistent with zero. Any statistically significant deviation from zero at any values of $Y^{2}$ would therefore suggest new physics beyond the SM, such as the existence of sterile neutrinos. Note that $Y^{2}$ could be experimentally accessed through the four-momentum different of measured final state lepton momentum and the known initial four-momentum. All other neutrino-related degrees of freedom that are not experimentally accessible have been fully integrated out at the theoretical level. As a result, the predicted polarization-dependent observables are expressed entirely in terms of experimentally measurable kinematics.